OpenStudy (anonymous):

prove the identity 1/(csc theta + cot theta) = csc theta - cot theta

OpenStudy (sepeario):

Use identities: csc(theta) = 1/cos(theta), cot(theta) = cos(theta)/sin(theta). If you have 1/cos(theta), then the cosine and sine reciprocate. You can combine fractions by finding the least common denominators. There's also the very important identity where: sin(theta)^2 + cos(theta)^2 = 1 Because this, you can subtract cosine^2 from both sides to get sin(theta)^2 = 1 - cos(theta)^2 Or the sine from both sides to get another identity. You can also divide both sides by cosine^2, to get: tan(theta)^2 + 1 = 1/cos(theta)^2 = sec(theta)^2 If you do the same with sine^2 you will get an identity involving csc and cot. Many identities. You must know them, or know how to derive them.

OpenStudy (anonymous):

I know identities.. struggling with algebra

OpenStudy (anonymous):

This is trigonometry...not algebra...